OpenStudy (anonymous):

x^2e^(-ax) How do I differentiate ?

OpenStudy (australopithecus):

split it into smaller functions that are easy to differentiate x^2 e^(-ax) then use product rule

OpenStudy (anonymous):

thank god, someone is here how do I know to use the chain rule or product rule first?

OpenStudy (anonymous):

how do I use product rule when there is that (-ax) part

OpenStudy (australopithecus):

Always split the function into smaller functions take the derivative then apply the rules

OpenStudy (australopithecus):

The derivative of e^u is (u')e^u

OpenStudy (australopithecus):

the derivative of e^x is (1)e^x = e^(x)

OpenStudy (anonymous):

how does that work into the product rule considering there is only f g f' g'

OpenStudy (australopithecus):

First make up two new functions out of this function so g(x) = x^2 s(x) = e^(-ax) find the derivative of both of these functions then just use product rule (MEMORIZE THIS) f'(x) = g'(x)s(x) + s'(x)g(x)

OpenStudy (australopithecus):

Their is really no trick to derivatives just split your function into the smallest functions possible then just apply rules

OpenStudy (australopithecus):

you really only need to memorize a few chain, product rule, quotient rule (although you can just use product rule instead)

OpenStudy (anonymous):

I know the product rule I guess my question is this if I had X^2e^x I would just use product rule cause its two things X^2 and e^x, one is f the other is g But with e^-ax there is also the derivative of -ax, where does that fit in, into the product rule?

OpenStudy (australopithecus):

is that \[e^{-ax}\] or \[e^{-x} a\]

OpenStudy (anonymous):

the first one

OpenStudy (australopithecus):

have you covered partial derivatives in class?

OpenStudy (australopithecus):

? alright

OpenStudy (anonymous):

Yes I think so sorry I was looking at it here

OpenStudy (anonymous):

im not sure what a partial derivative is though

OpenStudy (australopithecus):

well I would just treat it as a constant

OpenStudy (australopithecus):

Do you know how to compute the derivative or do you still need help?

OpenStudy (anonymous):

oh boy yea I would like to know once and for all what the derivative is :)

OpenStudy (australopithecus):

well split it into smaller functions g(x) = x^2 s(x) = e^(-ax) finding the derivative of e^(-ax), split it further m(x) = e^(x) d(x) = -ax apply (d'(x))m(dx) then just put it all together and apply product rule

OpenStudy (australopithecus):

d(x) = -ax d'(x) = -a(1) = -a

OpenStudy (australopithecus):

s(x) = e^(-ax) s'(x) = (-a)e^(-ax) g(x) = x^2 g'(x) = 2x f'(x) = g(x)s'(x) + g'(x)s(x)

OpenStudy (anonymous):

ok so s'(x) is both the derivative of (-ax) and e^(-ax) this is what I was not sure of

OpenStudy (australopithecus):

yeah

OpenStudy (australopithecus):

the derivative of e^f(x) is f'(x)e^f(x)

OpenStudy (anonymous):

yes I am familiar with that

OpenStudy (australopithecus):

so e^(-ax) assuming, "a" is a constant would just be the same

OpenStudy (australopithecus):

if it was a third variable you would have to use partial derivatives

OpenStudy (anonymous):

the answer is 0, and 2/a

OpenStudy (australopithecus):

f'(x) = 2x(e^(-ax)) + x^2(-ae^(-ax))

OpenStudy (anonymous):

im trying to figure that out now so I get 2x(e^(-ax)+x^2*-ae^(-ax)

OpenStudy (australopithecus):

so the given answer is 0 or 2/a?

OpenStudy (anonymous):

sweeet

OpenStudy (anonymous):

x= 0, 2/a so both

OpenStudy (australopithecus):

Are you looking to make f'(x) = 0

OpenStudy (anonymous):

find a formula for the x coordinates fo the critical points of f in terms of a

OpenStudy (australopithecus):

alright fair enough

OpenStudy (australopithecus):

I'm thinking if a is your y value then you need to use implicit differentiation

OpenStudy (anonymous):

how do you simplify the equation we got Im having a hard time getting it down to 2/a

OpenStudy (australopithecus):

critical points are where your f'(x) = 0

OpenStudy (anonymous):

no this is just introductory derivatives from calc one before L Hopitals rule and everything this is just after learning chain rule, product rule

OpenStudy (australopithecus):

ok, but yeah critical points are found by setting f'(x) = 0 and also by checking the domain of f'(x)

OpenStudy (australopithecus):

critical points are an indication where the f(x) changes

OpenStudy (anonymous):

yea I reduce and then set to zero

OpenStudy (anonymous):

yea crit points are f'(x) inflection points are f''(x) (don't need that for this equation though)

OpenStudy (anonymous):

so you factor out e^(-ax)?

OpenStudy (australopithecus):

0 = 2x(e^(-ax)+x^2*-ae^(-ax) so one of the critical points will be 0 0 = 2(0)(e^(-ax)+(0)^2*-ae^(-ax) 0 = 0 I dont see any other critical points possible

OpenStudy (australopithecus):

So you haven't seen implicit differentiation?

OpenStudy (anonymous):

nope we don't deal with 3 variables yet

OpenStudy (australopithecus):

implicit differentiation isn't dealing with three variables

OpenStudy (australopithecus):

we are taking the derivative da/dx I think

OpenStudy (anonymous):

well in this example I have to algebraically simplify then solve both sides to 0

OpenStudy (australopithecus):

0 = xe^(-ax)(2-ax)

OpenStudy (australopithecus):

On implicit differentiation, if it is then we have to do this derivative differently https://www.math.ucdavis.edu/~kouba/CalcOneDIRECTORY/implicitdiffdirectory/ImplicitDiff.html

OpenStudy (anonymous):

ah ha! yes that last equation is it!

OpenStudy (australopithecus):

oh I guess a is a constant

OpenStudy (australopithecus):

I'm just tired ha

OpenStudy (australopithecus):

yeah I just factored it

OpenStudy (anonymous):

I set xe^(-ax) to 0 and get 0 then I solve for x in 2-ax to get 2/a?

OpenStudy (australopithecus):

set (2-ax) = 0 set xe^(-ax) = 0 and solve for x

OpenStudy (australopithecus):

remember e^(x) =/= 0

OpenStudy (anonymous):

I have a question can 2-ax=0 be both 2/a and a/2?

OpenStudy (australopithecus):

plug them in and see

OpenStudy (australopithecus):

show me what you get if you are still confused

OpenStudy (australopithecus):

you should help highschool students here with their algebra and use wolfram alpha to prove stuff to yourself. Wolframalpha.com can do calculus for you to btw

OpenStudy (australopithecus):

2/a and a/2? 2-a(2/a) = 0 \[ 2-a^1 2a^{-1} = 0\] add exponents of a 1 + (-1) = 1 - 1 = 0 \[ 2-a^0 2 = 0\] \[ 2-(1)2 = 0\] remember anything with the exponent 0 equals 1, excluding 0^0 which is undefined as for the other thing you suggested \[ 2-2\frac{a}{2} = 0\] \[ 2- a^1a^1 = 0\] \[ 2- a^2 = 0\] we dont know if a = (2)^(1/2) is so yeah it isn't a definite critical point

OpenStudy (australopithecus):

well anyways I'm leaving hope I was helpful if this was implicit differentiation check that link out, if it is calc 1 you will have to learn it anyways so might be worth checking out, you might also want to check khan academy or patrickmjt or what ever his name is on youtube

OpenStudy (anonymous):

thank you sooooo much